### v4

# Training Hybrid Deep Quantum Neural Network for Efficient Reinforcement Learning

Abstract
Quantum circuits embed data in a Hilbert space whose dimensionality grows exponentially with the number of qubits, allowing even shallow parameterised quantum circuits (PQCs) to represent highly–correlated probability distributions that are costly for classical networks to capture. Reinforcement-learning (RL) agents, which must reason over long-horizon, continuous-control tasks, stand to benefit from this expressive quantum feature space, but only if the quantum layers can be trained jointly with the surrounding deep-neural components. Current gradient-estimation techniques (e.g., parameter-shift rule) make such hybrid training impractical for realistic RL workloads, because every gradient step requires a prohibitive number of circuit evaluations and thus erodes the potential quantum advantage. We introduce qtDNN, a tangential surrogate that locally approximates a PQC with a small differentiable network trained on-the-fly from the same minibatch. Embedding qtDNN inside the computation graph yields scalable batch gradients while keeping the original quantum layer for inference. Building on qtDNN we design hDQNN-TD3, a hybrid deep quantum neural network for continuous-control reinforcement learning based on the TD3 architecture. On the high-dimensional Humanoid-v4 benchmark, our agent reaches a test return that surpasses classical TD3, SAC and PPO baselines trained with identical compute. To our knowledge this is the first PQC-enhanced policy that matches or exceeds state-of-the-art classical performance on Humanoid. qtDNN has the potential to reduce quantum-hardware calls significantly and is designed to be compatible with today’s NISQ devices. The method opens a path toward applying hybrid quantum models to large-scale RL and other gradient-intensive machine-learning tasks.
1 Introduction
Quantum computing has initially attracted significant attention for its potential to solve classically challenging problems by exploiting data representations in a Hilbert space that grows exponentially with the number of qubits, as discussed by Biamonte et al. (2017). Machine learning (ML) is an application field that could benefit greatly from quantum computing, as patterns are learned from increasingly large datasets (cf.  Vaswani et al. (2023); DeepSeek-AI et al. (2025b)) and more efficient methods are necessary to advance artificial intelligence (cf. Robison (2024); Samsi et al. (2023); Wells (2025)). Quantum machine learning (QML) aims at enhancing the efficiency and generalizability of learning beyond classical methods (see Table 1 in Biamonte et al. (2017) for known speedups of various subroutines). Although fault-tolerant quantum machines capable of running complex circuits without error are not yet available to exponentially accelerate computing with algorithms based on Harrow et al. (2009) or Brassard et al. (2000), the rapid development of noisy intermediate-scale quantum (NISQ) devices has inspired a new class of QML techniques designed to operate under noisy constraints (cf. Wang and Liu (2024)). On one hand, recent quantum supremacy experiments demonstrated that NISQ quantum devices can generate distributions where measured qubit states exhibit correlations that can not be simulated with reasonable classical compute (cf. Arute et al. (2019); Acharya et al. (2025); Gao et al. (2025)). On the other hand, a subfield of ML, reinforcement learning (RL) has become a central pillar of modern AI (cf. ACM (2025)), by estimating classical probability distributions that capture regularities in its interaction data (cf. Vaswani et al. (2023)), which allows an RL agent to refine its policy to attain high performance on target tasks (cf. Schulman et al. (2017); Haarnoja et al. (2018); DeepSeek-AI et al. (2025a)). Nontrivial highly correlated distributions seen in sampled output data (cf. Andersen et al. (2025)) from NISQ QPUs have the potential to better model complicated correlations in applications (cf. Lo et al. (2023); Huang et al. (2022); Vaswani et al. (2023)), allowing RL models to achieve higher rewards, data efficiency, and robustness. Recent work from Jin et al. (2025) has shown that SOTA quantum reinforcement learning (QRL) with PPO modified to incorporate parametrized quantum circuits (PQCs), could have similar performance as the classical baseline models from Raffin (2020) on simple reinforcement learning benchmarks in Towers et al. (2024). This hybrid quantum machine learning is realized as a hybrid deep quantum neural network (hDQNN). In this setup, a classical deep neural network first maps the input into a vector of real numbers, those numbers become the control parameters fed to the downstream PQC during the forward pass. The input data to the PQC is encoded into the Hilbert space spanned by the $N$ qubits, $\mathcal{H}_{N\text{q}}\cong\mathbb{C}^{2^{N}}$. Low-dimensional classical input and projected (sampled) output act as bottleneck layers. QRL has shown a lot of promise when it comes to tackling tricky reinforcement learning tasks, especially those involving high-dimensional state and action (discrete and continuous) spaces. (cf. Dong et al. (2008); Paparo et al. (2014); Dunjko et al. (2017); Jerbi et al. (2021); Wu et al. (2025)) However, a significant barrier for applying hDQNNs to such high-dimensional problems is the cost of back-propagation through the PQC to update weights of classical DNNs before it.
Although methods exist for estimating gradients of a given PQC (cf. Wang and Liu (2024)), these can be impractical for real quantum hardware when the input and output dimensions are large as they require a large number of sequential quantum circuit measurements and are further incompatible with mini-batch updates due to the lack of parallel processing in physical QPUs. The lack of high-throughput training strategies compatible with NISQ hardware limits QRL’s practical performance and applicability to real-world tasks.
In this work, we introduce hDQNN-TD3, a hybrid quantum neural network architecture that solves this bottleneck using a novel approach: a classical surrogate that we named qtDNN will approximate the PQC’s gradient updates on the backward pass. We demonstrate the hDQNN-TD3 model on the Humanoid-v4 reinforcement learning benchmark with better performance over the widely adopted classical SOTA counterparts. This paper is organized as follows: we will introduce the method that enables scalable training of hDQNN with classical DNNs around PQCs in Sec. 2. Following this, Sec. 3 will introduce hDQNN-TD3 for continuous variable RL tasks and show the experimental results from Humanoid-v4 demonstrating that hDQNN-TD3 outperforms comparable classical methods as well as publicly reported models using SOTA RL architectures from Zhao (2024, 2023a, 2023b). The last Sec. 4 will discuss outlooks to train significantly larger models.
2 Neural PQC Tangential Approximation
2.1 Back-propagation and the quantum bottleneck
Modern machine learning methods are established with highly efficient backpropagation-based training: each operation in the computation graph supplies analytical Jacobians, so the loss can be differentiated w.r.t. millions of parameters and "backpropagated" through them with a single reverse pass. The updated parameters are used for a new forward pass, which creates a new differentiable loss. This process repeats until the loss reaches a set convergence condition.
On the other hand, hDQNNs are typically developed using PQCs to act as functional layers (QL) without learnable parameters. The described flow is typically interrupted by a QL  [^1] QL: produces output from the expectation or most probable outcome over $S$ shots of PQC, parametrized by its input, executed on simulator or physical QPU.. Fig. 1a shows the typical layout: a classical neural network PreDNN produces PQC control parameters $\mathbf{q}_{\text{i}}$, a QL outputting a bitstring $\mathbf{q}_{\text{o}}$ after executing PQC $S$ times, and a second classical neural network PostDNN maps $\mathbf{q}_{\text{o}}$ to the final prediction. $d_{\text{c-link}}$ classical links directly connecting PreDNN and PostDNN are added on top of the QL. These direct links allow classical communication between the DNNs around the QL, to help the PostDNN properly handle the QL’s sampled output.
It is however challenging to backpropagate loss efficiently through QLs. Unlike Gaussian re-parameterization layers, a PQC (and QL) has no closed-form gradient. Various strategies have been developed in recent years (cf.  Wang and Liu (2024)). Most methods obtain the gradient vector of a PQC with respect to its input parameters by varying each input parameter and using repeated PQC executions to estimate its partial derivative with respect to each varied input. Such methods become quickly intractable as it must run $\mathcal{O}(I)$ extra circuits per output dimension, driving the hardware cost to $I\times O\times S\times N_{b}$ evaluations per mini-batch for the QL with $I$ and $Q$ input and output dimensions respectively (cf. Sec. A.1.2). Despite requiring many repeated measurements that scale with the size of input and output, methods based on direct PQC gradient estimation make it intrinsically unfeasible to batch-backpropagate in parallel, prohibiting efficient mini-batched training for a hDQNN, especially when it is applied to complex reinforcement learning.
2.2 qtDNN: a tangential PQC surrogate trained on-the-fly
Figure 1: a, a typical hDQNN contains a PreDNN connected to a PostDNN via both $d_{\text{c-link}}$ direct connections and a quantum layer realized as a parametrized quantum circuit. The hDQNN inferences through the training dataset and records $(\mathbf{x},\mathbf{y}_{\text{pred}},\mathbf{y},\mathbf{q}_{\text{i}},\mathbf{q}_{\text{o}})$ of each forward pass into a buffer $\mathcal{M}$. In an update step, a batch, $\mathcal{B}$,of $N_{\text{b}}$ entries is sampled from $\mathcal{M}$ for updating hDQNN parameters. b, $\{(\mathbf{q}_{\text{i}},\mathbf{q}_{\text{o}})\}$ from $\mathcal{M}$ is stored into a qtDNN Replay Buffer. $N_{\text{qt}}$ tiny-batches, $\{\mathcal{B}_{\text{qt}}\}$, are sampled from $\mathcal{B}$ are then used to update qtDNN towards approximating PQC in $\mathcal{B}$. c, the updated qtDNN is then used in the surrogate model $Q_{\text{qt}}$ in this update step to facilitate the batched backpropagation of loss that are used to update the PreDNN and PostDNN with fixed qtDNN.
We resolve this bottleneck by introducing the qtDNN, a quantum tangential space learning architecture that approximates the local behaviors of a QL with a classical DNN. The qtDNN is a small, differentiable surrogate network that locally mimics the PQC inside each mini-batch. It serves as a component of the hDQNN that substitutes the PQC whenever differentiability is needed.
The qtDNN is used during the training loop to provide a differentiable first-order approximation of PQC’s behaviors, local to each minibatch during the backpropagation. Meanwhile, the forward inference pass goes through the QLs. To make training the qtDNN more efficient, we add a replay buffer (qtDNN-buffer) to each QL that stores the input, $\mathbf{q}_{\text{i}}$, and sampled output, $\mathbf{q}_{\text{o}}$, of the quantum layer as a pair, $(\mathbf{q}_{\text{i}},\mathbf{q}_{\text{o}})$, from each minibatch. This qtDNN-buffer is then used to train the qtDNN repeatedly with $N_{\text{qt}}$ tiny-batches randomly sampled from this qtDNN-buffer. This accelerates qtDNN’s convergence to a moving approximation to its target quantum layer (cf. Haarnoja et al. (2018)).
The update epoch is split in two stages (Fig. 1b–c):
  1. 1.
    Surrogate fit: From the sampled mini-batch =$\mathcal{B}=\{(\mathbf{q}_{\text{i}},\mathbf{q}_{\text{o}})\}$ we draw $N_{\mathrm{qt}}$ tiny-batches to minimize a binary-cross-entropy between $\text{qtDNN}(\mathbf{q}_{\text{i}})$ and the observed bitstrings $\mathbf{q}_{\text{o}}$.
  2. 2.
    Network update: We freeze the qtDNN, replace the PQC by it in a surrogate model $Q_{\text{qt}}$ that is otherwise identical to hDQNN, and run ordinary batched back-propagation to update the classical weights of the hDQNN.
Because the qtDNN is itself a conventional MLP, gradients propagate without further quantum calls. The next subsection formalizes the approximation guarantee.
2.3 Local gradient fidelity of qtDNN
Intuitively, the PQC is a smooth function of its parameters over the small region explored by one mini-batch. A feed-forward MLP can capture its first-order behavior. Theorem 1 states this precisely:
Theorem 1 (Local gradient fidelity of qtDNN).
Let $Q:\mathbb{R}^{d}\to\mathbb{R}^{k}$ be the deterministic function realised by a PQC under the expectation over $S$ shots, and suppose $Q$ is continuously differentiable on an open set $\mathcal{U}\subset\mathbb{R}^{d}$. Fix a mini-batch =$\mathcal{B}=\{x_{i}\}_{i=1}^{N_{b}}\subset\mathcal{U}$ lying in a closed ball of radius $r$ centered at $\bar{x}=\frac{1}{N_{b}}\sum_{i}x_{i}$. For every pair of tolerances $(\varepsilon_{1},\varepsilon_{2})>0$ there exists a feed-forward MLP $Q_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{k}$ such that:
$\max_{x\in\mathcal{B}}\bigl{\|}Q_{\theta}(x)-Q(x)\bigr{\|}\leq\varepsilon_{1},\qquad\max_{x\in\mathcal{B}}\!\bigl{\|}\nabla Q_{\theta}(x)-\nabla Q(x)\bigr{\|}\leq\varepsilon_{2}.$ (1)
Consequently, if the PQC is replaced by $Q_{\theta}$ inside the computation graph during back-propagation, the mini-batch gradient with respect to any upstream parameter vector $w$ is perturbed by at most an additive $O\!\bigl{(}\max(\varepsilon_{1},\varepsilon_{2})\bigr{)}$ term.
The first inequality is simply the qtDNN learning to approximate the PQC’s outputs correctly, while the second one shows that it learns to approximate the gradient of the PQC had it been differentiable i.e. approximate its effect during the backpropagation.
Intuition behind the proof.
Because the mini-batch $\mathcal{B}$ lies inside a small radius-$r$ ball, $Q$ is well-approximated by its first-order Taylor expansion around the centre $\bar{x}$. Thus, we (i) approximate the first-order Taylor polynomial $P(x){:=}Q(\bar{x}){+}J_{Q}(\bar{x})(x{-}\bar{x})$ over $\mathcal{B}$ by an MLP $Q_{\theta}$, and (ii) use Lipschitz bounds on the second derivative of $Q$ to control the Taylor remainder and its gradient. Replacing the PQC by that MLP therefore changes (i) forward activations by at most $\varepsilon_{1}$ and (ii) their Jacobian by at most $\varepsilon_{2}$, which in turn perturbs the gradient flowing into earlier layers by the same order of magnitude. See Appendix A.2 for full details.
3 Quantum Reinforcement Learning
3.1 Problem set-up and motivation
We now instantiate the qtDNN surrogate inside a continuous–control agent (Fig. 2a). The target task is Humanoid-v4 $(d_{\text{obs}}=376,d_{\text{act}}=17$), a long–horizon MuJoCo benchmark considered challenging for both classical and quantum methods ( Towers et al. (2024)). Classical baselines trained with identical compute on public HuggingFace repositories obtain final episode returns PPO $\approx 685$, SAC 5 742absent5742$\approx$ 5\,742≈ 5 742, TD3 5 306absent5306$\approx$ 5\,306≈ 5 306Zhao (2024),  Zhao (2023a),  Zhao (2023b)). We also saw that the best seed of our hybrid agent surpasses all three with 6 01160116\,0116 011 (see §3.5.1).
Figure 2: a, the typical exploration flow for the Agent to interact with the reinforcement learning objective’s target environment, $\mathcal{E}$, and obtain experience data to be stored in the replay buffer, $\mathcal{M}$. b, the particular modular model architecture used for the Actor Model, $A$, that allows the intermediate mapping $T$ that maps its input vector $\mathbf{q}_{\text{i}}$ to $\mathbf{q}_{\text{o}}$ to be implemented differently without changing the rest of the model design. This allows comparing learning performance of different quantum and classical implementations fairly.
Actor design. The actor consists of a classical pre-network PreDNN, a middle block $T$, and a post-network PostDNN (Fig. 2b). We instantiate $T$ as:
  1. 1.
    PQC (our hDQNN-TD3 model),
  2. 2.
    fully–connected (FC),
  3. 3.
    random bit generator (RBG), or
  4. 4.
    all-zero layer (0L)
to create a controlled ablation.
3.2 hDQNN-TD3 overview
We start from Twin-Delayed DDPG (TD3) because it is competitive yet simple. As the Actor Model is what drives the actions, a more generalizable Actor Model will directly impact the performance. Therefore, the original TD3 critic remains classical, only the actor contains the PQC (or its surrogate). The qtDNN is trained inside each mini-batch to approximate the PQC locally, enabling standard batched back-propagation through the whole actor. We call the resulting model hDQNN-TD3. Fig. 2 sketches the interaction loop between the Agent and the classical environment $\mathcal{E}$.
For clarity we recap the four alternating phases that occur during training, the symbols follow the notation introduced in Subsec. 2.2.
1. Exploration.
At every environment step the current state $\mathbf{s}\in\mathcal{S}$ is encoded by the PreDNN. Its output is split into (i) classical bypass features $d_{\text{c-link}}$, and (ii) the control vector $\mathbf{q}_{\mathrm{i}}$ that parameterises the PQC layer. Sampling the PQC yields the binary vector $\mathbf{q}_{\mathrm{o}}$, which the PostDNN transforms—together with the bypass features—into the continuous action $\mathbf{a}\in\mathcal{A}$. Applying $\mathbf{a}$ in $\mathcal{E}$ returns the next state $\mathbf{s}^{\prime}$, reward $r$, and a done flag. The transition tuple $(\mathbf{s},\mathbf{a},\mathbf{q}_{\mathrm{i}},\mathbf{q}_{\mathrm{o}},r,\mathbf{s}^{\prime})$ is appended to the replay buffer $\mathcal{M}$.
2. qtDNN update.
After $\mathcal{M}$ holds at least $N_{b}$ transitions we sample a mini-batch $\mathcal{B}\subset\mathcal{M}$. For every PQC in the Actor we collect the pairs $(\mathbf{q}_{\mathrm{i}},\mathbf{q}_{\mathrm{o}})$ of the batch and train its qtDNN surrogate with $N_{\mathrm{qt}}$ tiny-batches (from the same collection) using binary-cross-entropy. This freezes the qtDNNs for the remainder of the epoch.
3. Critic update (TD learning).
With qtDNNs fixed, the surrogate Actor $A_{\mathrm{qt}}$ provides target actions $\hat{\mathbf{a}}=A_{\mathrm{qt}}(\mathbf{s}^{\prime})$. The two critics minimize the TD-error $\epsilon_{\text{TD}}=\bigl{\|}r+\gamma\,C_{t}(\mathbf{s}^{\prime},\hat{\mathbf{a}})-C(\mathbf{s},\mathbf{a})\bigr{\|}^{2}$ as in TD3, using the loss averaged over $\mathcal{B}$.
4. Actor update.
Every $N_{\text{A}}$ critic steps we update the classical parameters of the Actor by maximizing $C(\mathbf{s},\,A_{\mathrm{qt}}(\mathbf{s}))$ through standard back-propagation that now flows through the differentiable qtDNN, thereby avoiding expensive parameter-shift evaluations of the PQC. Target networks $C_{t}$ and $A_{t}$ are softly updated with rate $\tau$.
Algorithm 1 exposes the training loop pseudo code for hDQNN-TD3.
Moreover, Fig. 5 in Appendix A.4 details the data flow inside one update epoch, making explicit which parameters are optimized at which stage.
Algorithm 1 hDQNN-TD3: TD3 actor with a PQC and qtDNN surrogate
1:Environment $\mathcal{E}$, replay buffer $\mathcal{M}$, actor $A$ (PQC inside), twin critics $C_{1},C_{2}$, target networks $A^{\prime}$, $C_{1}^{\prime},C_{2}^{\prime}$, update periods $N_{\text{A}}$, surrogate epochs $N_{\text{qt}}$
2:for episode $=1,2,\dots$ do
3:     collect experience: interact using $\mathbf{a}_{t}=A(\mathbf{s}_{t})+\mathcal{N}_{t}$, store $(\mathbf{s}_{t},\mathbf{a}_{t},r_{t},\mathbf{s}_{t+1})$ in $\mathcal{M}$
4:     if time to update then
5:         Sample mini-batch $\mathcal{B}$ of size $N_{\text{b}}$ from $\mathcal{M}$
6:         (a) update qtDNN: train surrogate $A_{\text{qt}}$ for $N_{\text{qt}}$ tiny-batches $\{(\mathbf{q}_{\text{i}},\mathbf{q}_{\text{o}})\}\in\mathcal{B}$
7:         (b) critic step: TD target $y=r+\gamma\min_{j=1,2}C_{j}^{\prime}\bigl{(}\mathbf{s}^{\prime},A^{\prime}(\mathbf{s}^{\prime})\bigr{)}$; regress $C_{1},C_{2}$ to $y$
8:         if global step mod $N_{\text{A}}=0$ then
9:              (c) actor step: maximize $C_{1}(\mathbf{s},A_{\text{qt}}(\mathbf{s}))$ w.r.t. classical actor parameters
10:              (d) target update: $\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime}$
11:         end if
12:     end if
13:end for
3.3 Quantum–classical instantiation details
The actor’s middle block $T$ is a 10-qubit, 10-layer PQC. Each layer applies parameterized single-qubit rotations followed by at most one CZ entangler, and all controls come from the preceding PreDNN. Concretely:
  • The input data is encoded in the controlling parameters $\{\theta_{ij},\phi_{ij},p_{j},k_{j}\}$.
  • $N=10$ qubits, initialized in |$\ket{0}^{\otimes N}$.
  • $M=10$ successive layers of gates.
  • Single-qubit stage. For layer $j$ and qubit $i$ we apply $\mathbf{R}\bigl{(}\theta_{ij},\phi_{ij}\bigr{)}$, where $\theta_{ij},\phi_{ij}\in[0,2\pi)$ are two outputs of PreDNN. Across the layer this uses $2N$ neurons.
  • Entangling stage. The same layer chooses a control–target pair $(p_{j},k_{j})\in\{0,\dots,N-1\}^{2}$ by rounding two additional PreDNN outputs and applies a CZ gate if $p_{j}\neq k_{j}$ (identity otherwise).
  • Hence PreDNN provides $(2N+2)M=220$ real controls per actor forward pass.
All remaining architectural details222mini-batch size $N_{\text{b}}=2^{8}$,actor update frequency $N_{\text{A}}=2$., including qtDNN width ($L_{\text{h}}=2\,048$), critic topology, learning rates, and exploration noise, are listed in Table 1, where $\alpha$ is the learning rate and the Critic follows the structure from Lillicrap et al. (2019).
Table 1: hDQNN Reinforcement Learning Model Configuration for Humanoid-v4 Problem
DNN Design Hyperparameters
Components Layers Activation $N$,$M$ $d_{\text{c-link}}$ $\alpha$ $\gamma$ $\tau$
PreDNN [376,256,230] Leaky ReLU 10 $3e^{-4}$ $5e^{-3}$
qtDNN [220,2 048,10] Leaky ReLU $3e^{-4}$
PQC 10,10
PostDNN [20,256,17] Leaky ReLU 10 $3e^{-4}$ $5e^{-3}$
Critic [[376,256],[273,256,1]] Leaky ReLU $3e^{-4}$ 0.99 $5e^{-3}$
3.4 Experimental protocol
We train each agent for up to 7 20072007\,2007 200 episodes (750 000absent750000$\approx$ 750\,000≈ 750 000 timesteps) across four seeds. [^3] The horizon of $7,200$ episodes balances limited compute resources against the large number of ablation variants we evaluate. A single longer-horizon run is provided in §3.5.1.
The qtDNN uses a single hidden layer of size $L_{\text{h}}=2\,048$ and $N_{\text{qt}}=32$ updates per actor step. Other parameters are exposed in Table 1. PPO/SAC/TD3 baselines follow the best Hugging Face RL models on official Gymnasium Humanoid-v4 benchmark and consume the same number of steps to ensure a fair comparison.
Ablation.
For the ablation in §3.5.2 we keep the same total number of environment episodes and we note that there is an error rate $\sim 10^{-3}$ associated with each PQC gate. Within the same architecture, we run the experiment for an FC layer, an RBG layer, and the PQC layer across $n_{\text{qubits}}\in\{5,10\}$ and $n_{\text{shots}}\in\{100,1\,000\}$. We omit the $(n_{\text{qubits}},n_{\text{shots}})=(5\,100)$ configuration because the run got impractically slow for some seeds (see Appendix A.1 for more details on limitations). A single-qubit setting would behave almost classically, whereas circuits with more than $\approx 20$ qubits can trigger peaks above 30 GB in GPU memory, leaving too little head-room on the 40 GB GPU (cf. §3, Google Quantum AI (2025)). Likewise, $10$ shots are statistically insufficient and 10 0001000010\,00010 000 shots yield negligible variance reduction relative to their cost.
3.5 Results
3.5.1 General performance
In the following part, we will observe the mean test return across training sessions. The test return is recorded by averaging episode rewards of 10 test episodes every 10 000 training steps (approximately 10 training episodes when the agent can constantly finish the whole 1 000 steps of an episode). The graphs show a 10-episode rolling average.
(a) Mean test return (10-episode rolling average).
(b) Policy loss.
Figure 3: Learning dynamics of the hDQNN agent on Humanoid-v4. Shaded bands denote $95\%$ confidence intervals across seeds.
Learning dynamics across seeds.
Fig. 3 aggregates the four independent training runs across 7 20072007\,2007 200 episodes and reveals three salient trends. First, the mean test return (Fig. 3(a)) rises almost monotonically after the initial warm-up, surpassing the $10^{3}$ mark around episode 3 00030003\,0003 000 and accelerating sharply once the actor has experienced 5 50055005\,5005 500 episodes. The shaded $95\%$confidence band widens in this late phase, indicating that one or two seeds surge ahead while the slower runs continue to improve, albeit more gradually. Second, the policy loss (Fig. 3(b)) mirrors this behavior in reverse: it drops rapidly to $-100$ during the first 1 00010001\,0001 000 episodes, then decays almost linearly to $-200$ before steepening again in tandem with the return surge. The broader confidence interval after episode 5 50055005\,5005 500 signals larger gradient magnitudes and increased exploration of high-return regions. Finally, the concurrent monotonic gain in return and decline in loss across all seeds confirms that the qtDNN surrogate supplies gradients of sufficient fidelity throughout training, while the late-stage divergence illustrates the head-room the quantum-enhanced actor can exploit once it enters the high-performance regime.
Figure 4: Mean test return (10-rolling average of each actor’s best seed).
Best-seed comparison with benchmark.
Fig. 4 contrasts the best seed of each actor variant over a longer 14 0001400014\,00014 000-episode horizon. [^4] For clarity the PPO, SAC and TD3 reference values from best Hugging Face RL models (Zhao (2024, 2023a, 2023b)) on Gymnasium Humanoid-v4 are shown as horizontal dashed lines. The hDQNN-TD3 actor (solid blue) climbs steeply after episode 5 50055005\,5005 500 and saturates near a 6 011±83plus-or-minus6011836\,011$\pm$ 836 011 ± 83 return, surpassing the best seeds of classical TD3 (5 30653065\,3065 306), SAC (5 74257425\,7425 742) and PPO ($685$). Two additional aspects are noteworthy:
  • Sample-efficiency. hDQNN-TD3 reaches the 3 00030003\,0003 000-return mark about $800$ episodes earlier than its fully connected (FC) alternative, suggesting that the quantum layer helps discover high-value policies sooner.
  • Stability of convergence. The hDQNN-TD3 curve flattens into a narrow band after episode 11 0001100011\,00011 000, whereas the FC and RBG variants continue to exhibit larger oscillations, this indicates that the qtDNN surrogate supplies gradients of sufficient fidelity to stabilize late-stage policy refinement.
Overall, the figure supports our claim that, in the best case, a single well-tuned hDQNN seed can match the ceiling performance of the strongest classical baselines while requiring fewer interaction steps to get there.
3.5.2 Ablation study
Table 2 contrasts five actor variants at two training checkpoints [^5] Since the test return is noisy, we used the test return’s 10-episode rolling average across episodes to smoothen it and avoid recording misleading noise at the checkpoint.. The upper block (Episode 4 00040004\,0004 000) gives a mid-training snapshot. Already at this stage the $10$-qubit, 1 00010001\,0001 000-shot PQC attains a mean return of $678.7\pm 179.3$, and its best seed reaches $926.7$, beating the FC by $+296$ (mean-to-mean) and the RBG by $+322$.
The lower block (Episode 7 20072007\,2007 200) captures late-training behavior. Here, the $5$-qubit, 1 00010001\,0001 000-shot PQC leads with a mean return of 1 132.6±1 411.4plus-or-minus1132.61411.41\,132.6$\pm$ 1\,411.41 132.6 ± 1 411.4 and a best seed of 3 211.63211.63\,211.63 211.6, maintaining a $+512$ margin over FC and $+706$ over RBG. The larger 10-qubit variant still improves on FC by $+149$ despite its higher circuit cost.
Replacing the PQC with either an RBG or a deterministic FC layer consistently lowers final performance by hundreds of return points, confirming that the quantum layer contributes non-trivial modeling capacity rather than acting as a mere regulariser.
Table 2: Ablation study on Humanoid-v4. Test returns are averaged over 4 seeds.
Episode Layer Shots Qubits Mean return Best return Time[min] PQC calls
4 00040004\,0004 000 RBG 0 0 $356.5\pm 219.1$ $545.3$ $39.8$ 0
4 00040004\,0004 000 FC 0 0 $382.7\pm 167.5$ $603.6$ $27.4$ 0
4 00040004\,0004 000 PQC $010$ $470.8\pm 57.1$ $534.5$ $291.2$ 302 738302738302\,738302 738
4 00040004\,0004 000 PQC $1000$ 5 $521.3\pm 128.0$ $650.8$ $227.3$ 302 948302948302\,948302 948
4 00040004\,0004 000 PQC $0010$ $678.7\pm 179.3$ $926.7$ $307.3$ 359 952359952359\,952359 952
7 20072007\,2007 200 RBG 0 0 $426.8\pm 114.5$ $545.0$ $105.9$ 0
7 20072007\,2007 200 FC 0 0 $620.8\pm 393.8$ 1 207.81207.81\,207.81 207.8 $89.8$ 0
7 20072007\,2007 200 PQC $010$ $517.8\pm 95.0$ $659.2$ $712.2$ 694 104694104694\,104694 104
7 20072007\,2007 200 PQC $1000$ 5 $1132.6\pm 1411.4$ 3 211.63211.63\,211.63 211.6 $737.9$ 769 867769867769\,867769 867
7 20072007\,2007 200 PQC $0010$ $768.9\pm 387.4$ 1 277.41277.41\,277.41 277.4 $700.9$ 761 498761498761\,498761 498
This consistent increase of performance between the FC and the PQC layer shows that the qtDNN is able to approximate the QL during back-propagation: the model is able to make use of the more complex representation of the QL during forward-prop and achieve a higher return, all while keeping stable gradients during back-propagation. Although not a direct experimental proof, this observation goes in the way of validating the local-tangent result of Theorem 1, i.e. $\|\nabla Q_{\theta}-\nabla Q\|\leq\varepsilon$ across the training.
3.5.3 Quantum calls
The PQC calls column from Table 2 shows how many quantum calls have been performed during the training. This number, $K$, is the same as the number of total steps the agent took during the training to explore its environment and populate the replay buffer, $\mathcal{M}$, for updating classical networks in off-policy training process. Each forward inference through the Actor Model makes one PQC call. Each PQC call to a quantum backend will return the estimated most probably bitstring from running identical PQC $S$ times (shots) and obtaining bitstrings from measuring all qubits in each shot. Therefore, hDQNN trained with qtDNN makes the minimal number ($K$, the same as the number of inference) of PQC calls and $S\times K$ PQC executions on an expensive physical QPU. Meanwhile, alternative quantum hardware compatible gradient methods like parameter shifts require $\sim I\times Q\times N_{\text{b}}\times S\times K$ additional physical QPU runs over the training process for our QL with $I=220$ dimensional input, $Q=10$ dimensional output, and a mini-batch size of $N_{\text{b}}=2^{8}$ in our cases.
4 Discussion
In this work, we proposed a novel way to efficiently train hDQNN compatible with NISQ quantum hardware, allowing for high-throughput mini-batch training using a surrogate classical neural network (qtDNN) to locally approximate and estimate gradients of the embedded quantum circuit for each mini-batch parameter update in off-policy reinforcement learning process. This method only requires expensive PQC calls in inference steps and therefore minimizes the number of expensive PQC calls in training significantly while still allows for efficient parallel mini-batch back-propagation to update classical weights. We further constructed hDQNN-TD3 model to learn complex, high-dimensional, and continuous-variable reinforcement learning tasks like Humanoid-v4 successfully with a final reward score comparable to that of models with widely used SOTA architectures except the recently proposed MEow from Chao et al. (2024). We further carried out ablation study by swapping the PQC to alternative classical layers and observed consistently better performance with hDQNN-TD3 against all comparable classical alternatives.
Therefore, training hDQNN with qtDNN and inference with trained hDQNN on (noisy) quantum hardware offer a new route in reinforcement learning and AI research for addressing complicated real-world problems that are always stochastic, highly correlated, and even challenging to simulate classically. Further research in this direction could lead to novel models such as hDQNN-LLM where classical layers responsible for high-dimensional encoding and sampling in LLMs, such as DeepSeek-AI et al. (2025a), are replaced with QLs surrounded by classical DNNs to learn the strongly correlated distributions in natural language as hinted by Lo et al. (2023). Moreover, the hDQNN-TD3 could be further developed to enhance models like NVIDIA (2025) in controlling advanced real-world robots.
Aknowledgements
This work has received partial funding from the European Union’s Horizon Europe research and innovation program under grant agreement No. 101084642. The work of G. K. has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program.
Disclaimer
This paper was prepared for information purposes and is not a product of HSBC Bank Plc. or its affiliates. Neither HSBC Bank Plc. nor any of its affiliates make any explicit or implied representation or warranty and none of them accept any liability in connection with this paper, including, but not limited to, the completeness, accuracy, reliability of information contained herein and the potential legal, compliance, tax or accounting effects thereof. Copyright HSBC Group 2025.
Appendix A Appendix
A.1 Limitations
A.1.1 Training hDQNN with qtDNN surrogate
The assumption that a quantum tangential deep neural network (qtDNN) can be trained as a good approximation to a given QL could only be reasonable for each mini-batch of training data (see Fig. 5) in off-policy reinforcement learning frameworks. It is because the sampled distribution of a general PQC is exponentially hard to simulate with classical computing resources Arute et al. [2019], Acharya et al. [2025].
Moreover, during the experiment, we implemented qtDNN as a 3-layer MLP and observed that training is most efficient when the hidden layer has $L_{\text{h}}=2^{N+1}$ neurons for a $N$-qubit wide PQC. As $N$-qubit spans a Hilbert space $\mathcal{H}_{N\text{q}}\cong\mathbb{C}^{2^{N}}\cong\mathbb{R}^{2^{N+1}}$, the experimentally observed best hidden layer’s linear space $\mathcal{L}_{\text{h}}\cong\mathbb{R}^{2^{N+1}}\cong\mathcal{H}_{N\text{q}}$ for $N=10$ in the experiment achieving a high score. Reducing the hidden linear space $\mathcal{L}_{\text{h}}$ to $\mathbb{R}^{2^{N}}$ has been observed to drastically reduce model performance in the final score. The observations indicate that physics-inspired appropriate multi-qubit state representation by the hidden layer is critical. Therefore, it is reasonable to expect, the number of neurons required for 3-layer MLP qtDNN scales according to $2^{N+1}$ with $N$ qubits in the PQC it approximates locally. We also note that the optimal 3-layer MLP design of qtDNN seems to be unexpectedly independent of PQC depth, $M$. This exponential resource scaling would limit the 3-layer MLP implementation of qtDNN to relatively narrow (in qubit count) PQC. The limitation on PQC width could further limit the ability for hDQNN + qtDNN to learn complex AI/ML tasks involving a vast number of correlated degrees of freedoms and hinder its applications in the Large Language Models (LLMs) DeepSeek-AI et al. [2025a, b] and advanced robotic controls in unpredictable real-world interactive conditions as in  NVIDIA [2025]. Future development of novel qtDNN architecture might significantly improve the scaling efficiency by tailoring the qtDNN design to respect the structure of a given PQC and bound its resource requirements with a function of $N_{\text{b}}$. For example, tensor network methods could approximate a PQC with a given structure efficiently Liu et al. [2024], Huang et al. [2021]. However, even with the potentially large resource requirements scaling exponentially with the number of qubits involved in sampling, we could argue that this limitation is only present in training and the inference data flow of a trained hDQNN is potentially exponentially more efficient as forward pass through the PQC could be carried out on physical QPUs.
For the inference, physical QPUs could carry out PQC operations directly on all qubits without the requirement of exponentially scaling physical computing resources Acharya et al. [2025]. Moreover, the superconducting QPU’s 10s ns level quantum entangling gate times Gao et al. [2025] are preferred for co-processing with classical neural networks because the gap in clock rates is much smaller as compared to trapped ion and neutral atom platforms’ average entangling gate times around 0.5 ms. Here, we note that even for SOTA high-connectivity superconducting quantum processors, each PQC call with $S=1,000$ will take about $t_{\text{qc}}\sim M\times 50~{}\text{ns}+100~{}\text{ns}\sim 0.6~{}\mu\text{s}$ (assuming a single-qubit and a two-qubit gates in sequence take 50 ns and measurement take 100 ns), leading to a per PQC Call to backend with $S=1,000$ shots consuming $S\times t_{\text{qc}}\sim 0.6~{}\text{ms}$ in inference step. Therefore, having fast gates is practically useful for quantum reinforcement learning tasks.
In terms of applying this hDQNN+qtDNN method to a variety of different RL environments, only Humanoid-v4 and the simple Pendulum-v1 were attempted due to time and resource limitations. Even though they represent two extremes of complexities in Gymnasium continous variable reinforcement learning benchmarks and achieved good results, the long PQC simulation time became a bottleneck making it challenging to extend the experimentation to other environments within the short time available to prepare the paper. However, it is straightforward to apply hDQNN + qtDNN for enhancing reinforcement learning in other useful scenarios involving both continuous and discrete observation and action spaces. Integration with the trailblazing AI models’ training and inference pipelines DeepSeek-AI et al. [2025a] could be among the topics in future researches.
In the end, hDQNN requires PQC to be executed on a physical or simulated QPU. Real-world QPUs in the near-term have noises (errors) in quantum operations they carry out. Therefore, in this work, we tested the noise tolerance of hDQNN + qtDNN method in learning Humanoid-v4. The hDQNN + qtDNN model having PQC bit-flip and phase-flip error rates of $10^{-3}$ performs on par and even potentially better than classical SOTA models in terms of rewards they achieved given a similar compute. However, a larger $10^{-2}$ error rate in physical quantum gates would lead to a final reward lower than SOTA models. On the other hand, we further noticed hDQNN + qtDNN model demonstrated faster learning with elevated error rates in quantum gates and attribute this to quantum gate noises (errors) having the same effect of larger exploration noises in classical SOTA models. We note that hyper parameters are kept the same between comparable alternative quantum and classical models presented in the main text. We further note that, as we took the most probable bitstring from sampling PQC over $S$ identical shots as the output $\mathbf{q}_{\text{o}}$ of a QL in hDQNN for a given input, $\mathbf{q}_{\text{i}}$, reducing the number of shots $S$, in each sampling for estimating $\mathbf{q}_{\text{o}}$ would increase the noises in $\mathbf{q}_{\text{o}}$. This increased sampling noise impacts the results in a similar way as errors in quantum operations.
A.1.2 The inefficiency of training hDQNN without qtDNN surrogate
Because $I\times O$ different circuits are required to estimate all the gradients of a QL with $I$ input dimensions and $O$ output dimensions and each circuit needs a number, $S$, of shots to have the statistics for estimating the observables used to calculate gradients. Moreover, learning for complex problems with a high sample data efficiency would require training data to be sampled from past experiences in mini-batches of size $N_{\text{b}}$ for each DNN update with batch optimization as in Haarnoja et al. [2018]. Therefore, existing QML methods that directly estimate gradients of the QL are faced with a prohibitive practical training overhead proportional to $I\times O\times S\times N_{\text{b}}$ for running on real QPU that can only run one shot at a time. For example, in the case of using a fast superconducting Josephson junction quantum processor that has gate time as short as 20 ns and measurement time in the range of 100 ns as seen in Acharya et al. [2025], Gao et al. [2025], a 10-layer 10-qubit PQC (QL) would need $I=220$ and $O=10$ with $N_{\text{b}}=256$ and $S=1000$ in our example for learning complex Humanoid-v4 environment. In this case, each update to model DNN parameters would require about 281.6 s. Considering that typical classical RL training even for simple tasks would likely require $\sim 10^{6}$ update steps similar to Haarnoja et al. [2018], this would translate to more than 10 years of training.
A.2 Proof of Theorem  1 on the Local gradient fidelity of qtDNN
Proof.
Let $\mathcal{B}\subset\mathbb{R}^{d}$ be the mini-batch, contained in the closed ball $\overline{B}_{r}(\bar{x})$. Because $Q$ is $C^{1}$ on an open neighbourhood of $\overline{B}_{r}(\bar{x})$, its Jacobian $J_{Q}$ is Lipschitz–continuous: we can fix a constant $L>0$ such that $\|J_{Q}(x)-J_{Q}(y)\|\leq L\|x-y\|$ for all $x,y\in\overline{B}_{r}(\bar{x})$. Set $M=\|\nabla^{2}Q(\bar{x})\|$ and $\delta=\max(\varepsilon_{1},\varepsilon_{2})-\frac{1}{2}Lr^{2}>0$.
Step 1. Approximating the affine Taylor part.
Write $h=x-\bar{x}$. First-order Taylor expansion gives $Q(x)=Q(\bar{x})+J_{Q}(\bar{x})h+\rho(x),$ where the remainder satisfies the integral form
[…]
Define the affine map $P(x):=Q(\bar{x})+J_{Q}(\bar{x})h$. The universal-approximation theorem ensures that there exists an MLP $Q_{\theta}$ such that
$\max_{x\in\mathcal{B}}\,\|Q_{\theta}(x)-P(x)\|\leq\delta,\qquad\max_{x\in\mathcal{B}}\,\|\nabla Q_{\theta}(x)-J_{Q}(\bar{x})\|\;\leq\;\delta.$
Step 2. Bounding the remainder $\rho$ and its gradient.
Because $J_{Q}$ is $L$-Lipschitz and $\|h\|\leq r$, from (A.2) we obtain
[…]
Differentiating (A.2) with respect to $x$ yields, using $\nabla_{x}h=I$,
[…]
Hence
[…]
Step 3. Uniform error bounds on $\mathcal{B}$.
For any $x\in\mathcal{B}$, combine (A.2) with (A.2):
$\|Q_{\theta}(x)-Q(x)\|\leq\underbrace{\|Q_{\theta}(x)-P(x)\|}_{\leq\delta}+\underbrace{\|\rho(x)\|}_{\leq\frac{1}{2}Lr^{2}}\leq\delta+\frac{1}{2}Lr^{2}=\varepsilon_{1}.$
Likewise, using (A.2) and (A.2),
$\|\nabla Q_{\theta}(x)-\nabla Q(x)\|\leq\underbrace{\|\nabla Q_{\theta}(x)-J_{Q}(\bar{x})\|}_{\leq\delta}+\underbrace{\|\nabla\rho(x)\|}_{\leq\frac{1}{2}(M+Lr)r}\leq\delta+\frac{1}{2}(M+Lr)r=\varepsilon_{2}.$
Step 4. Impact on back-propagation.
Inside the computation graph, replacing $Q$ by $Q_{\theta}$ changes the per-sample loss by at most $\varepsilon_{1}$, and perturbs every mini-batch gradient entering the upstream network by at most $\varepsilon_{2}$. Thus all parameter updates incur an additive $O\!\bigl{(}\max(\varepsilon_{1},\varepsilon_{2})\bigr{)}$ error, completing the proof.
A.3 Hardware and system software.
All numerical experiments were executed independently on a dedicated NVIDIA A100-SXM4 GPU (40 GB HBM2e). The software stack comprised:
  • Python 3.12.3;
  • PyTorch 2.6.0 built against CUDA 12.4.0;
  • Gymnasium 1.1.1 with MuJoCo 3.3.2;
  • cuda-q 0.9.1 for GPU-accelerated quantum-circuit simulation NVIDIA [2022].
A.4 Supplementary figures
Figure 5: The batched training process flow that incorporates qtDNN training framework for efficiently updating the classical neural layers of the hDQNN model to achieve desired learning goals.